# Identifying Spatiotemporal Patterns in Opioid Vulnerability: Investigating the Links Between Disability, Prescription Opioids and Opioid-Related Mortality

### v2

@@ Acknowledgments
The authors would like to thank Andrew Farrell for his helpful comments and feedback, not only on the manuscript, but throughout the entire study.
@@ 5.1   Author contributions
Appendix A Data uncertainty level sensitivity analysis
To choose a data uncertainty level for our study, we conducted a sensitivity analysis. This analysis examined various uncertainty levels, specifically 1%, 3%, and 5%, by evaluating the respective changes in the model’s predictive performance. Optimal predictive accuracy was found with a 1% data uncertainty level. Consequently, a 1% uncertainty assumption was adopted for all data measurements throughout our study.
Table 1 documents the affects of increasing the level of uncertainty in the data on the filter’s average general accuracy. The most significant drop in accuracy is observed in the disability rates dataset. The narrow range and tight clustering of these percentile ranks between 0 and 100 each year, combined with the use of synthetic data for intervening years, make this dataset particularly sensitive to increased uncertainty, due to a broader range of estimations from the filter.
Data Uncertainty Level Mortality Disability Dispensing
1% uncertainty 93.62% 95.35% 91.69%
3% uncertainty 93.45% 88.92% 90.03%
5% uncertainty 93.37% 88.64% 89.03%
Table 1: The affects on the filter’s general accuracy as the level of uncertainty in the data increases.
Appendix B The Kalman filter algorithm
@@
Appendix C Efficacy summary for the spatial Kalman filter during training years
Variable Year Avg General Acc Hotspot Acc Avg Error Max Error
Mortality rates 2015 96.58% 97.31% 0.31 9.07
2016 94.57% 96.55% 0.37 6.84
2017 94.21% 95.54% 0.40 6.98
2018 94.55% 94.34% 0.38 6.90
2019 93.87% 95.0% 0.38 6.14
Disability rates 2015 92.26% 99.38% 0.58 7.49
2016 91.75% 98.82% 0.77 9.33
2017 92.06% 99.41% 0.72 9.06
2018 92.73% 98.82% 0.86 11.80
2019 91.31% 97.59% 0.81 9.30
Dispensing rates 2015 96.47% 91.49% 0.46 13.15
2016 96.05% 87.36% 0.55 13.98
2017 96.53% 95.45% 0.46 13.26
2018 97.99% 91.86% 0.46 22.97
2019 97.50% 93.81% 1.81 72.28
Table 2: Efficacy metric results for the Kalman filter during the 2015 to 2019 training years
Appendix D Multi-year predictive efficacy of the spatial Kalman filter
@@
Figure 7: Histogram comparisons of the absolute errors for the 2020 (a) mortality, (b) disability, and (c) dispensing rate predictions between the fully trained filter and the filter trained on progressively less data. The error distribution of the fully trained filter is colored in blue, whereas the error distribution of the filter trained on less data is colored in red. The top row showcases comparisons for the filter trained on four years of data, iteratively progressing downwards to the last row showcasing comparisons for the filter trained on a single year of data. Each histogram features a red arrow highlighting the maximum error from the filter trained on less data.

### v3

@@ Acknowledgments
The authors would like to thank Andrew Farrell for his helpful comments and feedback, not only on the manuscript, but throughout the entire study. The authors would also like to thank the anonymous reviewers for their helpful comments and suggestions, which helped refine and strengthen the manuscript.
@@ 5.1   Author contributions
Appendix A The Kalman filter algorithm
@@
Appendix B Data uncertainty level sensitivity analysis
To choose a data uncertainty level for our study, we conducted a sensitivity analysis. This analysis examined various uncertainty levels, specifically 1%, 3%, and 5%, by evaluating the respective changes in the model’s 2020 predictive performance. Across all datasets, we observed a decline in accuracy as uncertainty increased. Given these findings, we opted to use a 1% uncertainty level for all data measurements throughout our study to maintain optimal predictive accuracy. Table 1 summarizes the impact of increasing data uncertainty on the filter’s 2020 average general accuracy.
Data Uncertainty Level Mortality Disability Dispensing
1% uncertainty 94.00% 88.49% 91.61%
3% uncertainty 93.87% 87.97% 89.96%
5% uncertainty 93.81% 87.70% 88.99%
Table 1: The affects on the filter’s 2020 general accuracy as the level of uncertainty in the data increases.
Appendix C Multi-year predictive efficacy of the spatial Kalman filter
@@
Figure 7: Histogram comparisons of the absolute errors for the 2020 (a) mortality, (b) disability, and (c) dispensing rate predictions between the fully trained filter and the filter trained on progressively less data. The error distribution of the fully trained filter is colored in blue, whereas the error distribution of the filter trained on less data is colored in red. The top row showcases comparisons for the filter trained on four years of data, iteratively progressing downwards to the last row showcasing comparisons for the filter trained on a single year of data. Each histogram features a red arrow highlighting the maximum error from the filter trained on less data. Figures are best viewed online in color.
Appendix D Impact of Initialization Year on Prediction Errors
In the previous section, we examined how reducing the Kalman filter’s training period while extending the forecasting horizon affected predictive errors. Here, we instead evaluate how shifting the initialization year affects the model’s 2020 predictions. Given the dynamic nature of opioid-related issues, the choice of historical data used to initialize the model may play a crucial role in capturing shifting trends. This analysis provides insight into the model’s adaptability to sudden changes in the data. If a significant shift occurs in opioid-related mortality, prescribing patterns, or social vulnerability factors, an extended training period may cause the model to overemphasize outdated trends rather than adjust to more recent dynamics. By iteratively shifting the initialization year forward, we assess whether relying on more recent observations enhances prediction accuracy or if a longer historical perspective remains beneficial.
Figure 8 presents the absolute error distributions for 2020 as a function of the initialization year, ranging from 2016 to 2019. In nearly all cases, the fully trained model—initialized with data from 2014 and trained through 2019—yields the more accurate results. However, there is one notable exception: When predicting 2020 opioid dispensing rates, the model initialized with 2019 data outperforms the fully trained model by a significant margin. As discussed earlier, major policy changes in 2019 led to a substantial drop in opioid prescriptions from 2018 to 2019. By contrast, from 2019 to 2020, dispensing rates remained stable without any dramatic shifts. The Kalman gain used in the fully trained model incorporates information about the sharp decline from 2018 to 2019, whereas the model initialized in 2019 does not, allowing it to better capture the stability of prescription rates from 2019 to 2020.
This result highlights an important consideration when selecting historical training data for time-series forecasting models: Abrupt shifts in a dataset can make older data less relevant—or even misleading—for future predictions. In stable data environments, a longer training period generally provides the best predictive performance by leveraging more comprehensive historical patterns. However, if a dataset contains abrupt changes from one time point to the next, due to policy shifts or other external factors, shorter training horizons may be preferable to avoid distortions caused by outdated trends.
Ultimately, our results suggest that shifting the initialization year can be a valuable strategy in contexts where sudden structural changes occur. Practitioners using Kalman filters or similar forecasting models should carefully weigh the trade-off between historical data depth and recent trend sensitivity. If the goal is to capture long-term trends, an extended training period is likely optimal. However, when significant, known disruptions exist, retraining the model on a more recent subset of data may improve short-term forecast accuracy.
Figure 8: Histogram comparisons of the absolute errors for the 2020 (a) mortality, (b) disability, and (c) dispensing rate predictions between the fully trained filter and the filter initialized at later years. The error distribution of the fully trained filter is colored in blue, whereas the error distribution of the filter initialized at later years is colored in orange. The top row showcases comparisons with the filter initialized at 2016, the second row for the filter initialized at 2017, the third row for the filter initialized at 2018, and the last row for the filter initialized at 2019. Each histogram features a blue arrow highlighting the maximum error from fully trained Kalman filter and an orange arrow highlighting the maximum error from the filter initialized at later years. Figures are best viewed online in color.
Appendix E Efficacy summary for the spatial Kalman filter during training years
Variable Year Avg General Acc Hotspot Acc Avg Error Max Error
Mortality rates 2015 96.57% 97.31% 0.31 9.13
2016 94.52% 97.04% 0.37 6.81
2017 94.25% 95.54% 0.41 7.07
2018 94.49% 94.34% 0.38 6.85
2019 93.83% 95.0% 0.38 6.15
2020 94.00% 70.00% 5.20 86.61
Disability rates 2015 92.25% 91.49% 0.58 7.49
2016 91.71% 89.66% 0.77 9.30
2017 92.04% 95.45% 0.72 9.07
2018 92.71% 91.86% 0.86 11.78
2019 91.31% 93.81% 0.81 9.30
2020 88.49% 94.19% 1.84 15.97
Dispensing rates 2015 96.47% 99.38% 0.46 13.16
2016 96.05% 98.82% 0.55 13.97
2017 96.53% 99.41% 0.46 13.26
2018 97.99% 99.41% 0.46 23.00
2019 97.50% 97.59% 1.81 72.32
2020 91.61% 68.64% 21.02 250.67
Table 2: Efficacy metric results for the spatial Kalman filter for all predictive years in study
Appendix F Data Summary Statistics
Variable Year Mean Std Dev Min Q1 Median Q3 Max
Mortality rates 2014 13.55 8.01 0.00 8.10 11.80 16.90 87.90
2015 14.84 8.83 0.00 8.90 12.90 18.50 113.90
2016 17.07 10.29 0.00 10.00 14.40 21.30 94.80
2017 18.46 11.42 0.00 10.80 15.60 23.00 141.50
2018 17.53 10.50 0.00 10.30 14.90 21.85 114.40
2019 18.01 10.78 0.00 10.70 15.40 22.30 117.30
2020 23.54 14.61 0.00 13.70 20.00 29.20 153.20
Disability rates 2014 49.64 28.94 0.00 24.45 48.93 74.59 99.97
2015 49.64 28.60 0.00 24.97 49.48 74.20 99.94
2016 49.63 28.95 0.00 24.62 49.38 74.50 100.00
2017 49.63 28.56 0.00 25.39 50.02 74.19 99.94
2018 49.63 28.96 0.00 24.83 49.06 74.34 100.00
2019 49.62 28.43 0.00 25.46 49.44 74.26 99.97
2020 49.62 28.96 0.00 24.47 49.49 74.63 100.00
Dispensing rates 2014 80.59 51.90 0.00 47.70 79.40 110.55 563.30
2015 75.10 47.88 0.00 45.60 73.50 103.30 504.90
2016 71.91 44.57 0.00 44.25 70.50 97.65 470.30
2017 64.61 40.13 0.00 39.50 62.70 87.55 402.00
2018 56.65 35.70 0.00 34.60 54.90 76.25 311.30
2019 40.49 33.11 0.00 19.55 34.60 54.00 567.90
2020 38.01 31.08 0.00 17.80 32.30 51.00 406.70
Table 3: Data summary statistics from 2014 to 2020.
Appendix G Comparison of the Spatial and Generic Kalman Filters
Figure 9: Histogram comparisons of the spatial and generic kalman filter’s absolute errors for the 2015, 2017, 2019 and 2020 (a) mortality, (b) disability, and (c) dispensing rate predictions. The error distributions of the spatial Kalman filter are colored in purple, whereas the generic Kalman filter’s error distributions are colored in orange. The top row showcases comparisons for year 2015, the second row for 2017, the third row for 2019, and the last row for 2020. Each histogram features a purple arrow highlighting the maximum error from spatial Kalman filter and an orange arrow highlighting the maximum error from the generic Kalman filter. Figures are best viewed online in color.